What is the width and length of a rectangle if the area and perimeter is 48 and 28
Answer by lwsshak3(11628)

What is the width and length of a rectangle if the area and perimeter is 48 and 28
***
let x=width
let y=length
..
xy=area=48
y=48/x
2x+2y=28
2x+2(48/x)=28
2x+96/x-28=0
lcd:x
2x^2+96-28x=0
x^2-14x+48=0
(x-8)(x-6)=0
x=8
or
x=6
width=6
length=8